OpenStudy (anonymous):

What does the notation \[ V(v_1, v_2)= f(v_1,v_2) (v_1+v_2) \] mean? What does the notation \[ V(v_1, v_2)= f(v_1,v_2) (v_1+v_2) \] mean? @Mathematics

OpenStudy (jamesj):

Context is important here but it looks like V is a function V : W x W --> W where W is a vector space and f is function f : W x W --> Reals (or whatever the scalars are here.)

OpenStudy (anonymous):

So what does it mean? Could you give an example?

OpenStudy (jamesj):

I can, but again, the context in which you find all of this is important as my hypothesis may not be right. Here's an example given my hypothesis. Let \[ f : \mathbb{R}^3 \times \mathbb{R}^3 \to \mathbb{R} \] be defined by \[ (v_1,v_2) \mapsto v_1 . v_2 \ \hbox{, the inner product/dot product} \] Then we could define a function \[ V : \mathbb{R}^3 \times \mathbb{R}^3 \to \mathbb{R}^3 \] by \[ V(v_1,v_2) = f(v_1,v_2) (v_1 + v_2) \]

OpenStudy (anonymous):

Hmm I think I misunderstood the actual text:) thanks for your help though :) I figured out what I did wrong as well:)